A battery with a terminal voltage of 12V is connected to a circuit consisting 

of four 10 Ω and one 5 Ω resistor all in series. Assume the battery has 

negligible internal resistance. Calculate the current through each resistor.


1
Expert's answer
2022-04-04T09:05:10-0400

Answer

Equivalent resistance is

"R=4*10+5=45\Omega"

These resistances are in series so current will be same.

"I=\frac{V}{R}\\=\frac{12}{45}\\=0.27A"